OpenStudy (anonymous):

i need answers for this to find out if it is right. 1. -6a + 3 = -3(2a-1) a. a=5 b. a=10 c. no solution d. identity .. 2. 3(c-2)=2(c-6) a. 2 b. -6 c. no solution d. identity .. 3. 0.5 (m+4) = 3(m-1) a. -2 b. 2 c.no solution d. identity thank you!!

OpenStudy (anonymous):

1. -6a + 3 = -3(2A-1)= A5 (A) 2. 3(c-2=2(c-6)=-6 due to C (B) 3. 0.5 (m+4) = 3(m-1) = Identity (D)

OpenStudy (anonymous):

thank you so much..

OpenStudy (anonymous):

@-StudyShark- check ur answer for first and third question

OpenStudy (anonymous):

first question answer is C and third question answer is B